OpenStudy (anonymous):

what is the solution of the linear equation x-2y+=-16 2x+y=18 somebody please

OpenStudy (anonymous):

First you have to get Y by itself. x-2y= -16 -2y= -16-x Divide -2 make sure its positive y y= 8+x y= x+8 either one. Second one goes the same 2x+y=18 -2x y=-2x+18

OpenStudy (anonymous):

x-2y+=-16 ------ (A) 2x+y=18 ------ (B) from (B) you get, by rearranging, y = 18 - 2x substitute this in (A) x - 2(18 - 2x) = -16 x - 36 + 4x = -16 5x = -16 + 36 x = 20/5 x = 8 substitute this in (B) 2(8) + y = 18 16 + y = 18 y = 18 - 16 y = 2 so solution is x = 8 and y = 2